Velma and Keota (V&K) is a partnership that owns a small company. It is considering two alternative investment opportunities. The first investment opportunity will have a five-year useful life, will cost $12,057.92, and will generate expected cash inflows of $3,100 per year. The second investment is expected to have a useful life of four years, will cost $9,513.82, and will generate expected cash inflows of $3,400 per year. Assume that V&K has the funds available to accept only one of the opportunities. (PV of $1 and PVA of $1) (Use appropriate factor(s) from the tables provided.)

Required

  1. Calculate the internal rate of return of each investment opportunity. (Do not round intermediate calculations.)

  2. Based on the internal rates of return, which opportunity should V&K select?

Internal Rate of Return
a. First investment %
Second investment %
b. V&K should select the
Answer #1

a)

First Investment:

IRR is the rate of return that makes initial investment equal to present value of cash inflows

12,057.92 = 3100 / (1 + R)1 + 3100 / (1 + R)4 + 3100 / (1 + R)3 + 3100 / (1 + R)4 + 3100 / (1 + R)5

Using trial and error method, i.e after trying various values for R, lets try R as 9.00%

12,057.92 = 3100 / (1 + 0.09)1 + 3100 / (1 + 0.09)4 + 3100 / (1 + 0.09)3 + 3100 / (1 + 0.09)4 + 3100 / (1 + 0.09)5

12,057.92 = 12,057.92

Therefore, IRR of First investment is 9.00%

Project y:

IRR is the rate of return that makes initial investment equal to present value of cash inflows

9,513.82 = 3400 / (1 + R)1 + 3400 / (1 + R)4 + 3400 / (1 + R)3 + 3400 / (1 + R)4

Using trial and error method, i.e after trying various values for R, lets try R as 16%

9,513.82 = 3400 / (1 + 0.16)1 + 3400 / (1 + 0.16)4 + 3400 / (1 + 0.16)3 + 3400 / (1 + 0.16)4

9,513.82 = 9,513.82

Therefore, IRR of Second investment is 16.00%

2)

Second investment should be accepted as it has the higher IRR